If you want to find the % increase from 281 students in 2008 to 1425 in 2009 how do you set this up to solve?
Answer by RAY100(1637)

% increase is the amt of increase divided by original number
.
(1425 - 281)/281 = 1144/281 =4.071 = 407.1 %
.